Find the equation of the line passing through the points (3,2) and (-3,8) Write the equation in point slope form, slope-intercept form, and standard form
Answer by ewatrrr(24785)

(3,2) and
(-3,8) m = -6/6 = -1
point(3,2) slope form: y - 2 = -1(x-3)
slope-intercept form: y = -x + 5
and standard form: x + y = 5